OpenStudy (anonymous):

how to simplify y^2+yw+w^2

OpenStudy (anonymous):

Depends on how you want it to look.. \[y^2 + yw + w^2\] \[=y^2 + yw + w^2 + 0\] \[=y^2 + yw + w^2 + (yw-yw)\] \[=y^2 + 2yw + w^2 -yw\] \[=(y+w)^2 - yw\] Is just one of many possible answers.